OpenStudy (anonymous):

solve the following DE by using the appropriate substitution: (y^2 + yx) dx - x^2 dy = 0

OpenStudy (anonymous):

( xy + y^2 ) dx + x^2 dy = 0 (xy + y^2 ) dx = - x^2 dy dy/dx = - ( xy + y^2 ) / x^2 On RHS, separating, dy/dx = - ( y/x) - (y/x)^2............(1) ......................................... Substitute y/x = v so that y = vx. Then, dy/dx = v + x ( dv/dx). ......................................... Hence, eq(1) now becomes v + x( dv/dx ) = - v - v^2 x( dv/dx ) = - 2v - v^2 = - v( 2 + v ) dv / v( 2 + v ) = - 2 dx / x Integrating both sides ( 1/2) INT [ 1/v - 1/ ( 2 + v ) ] dv = - 2 INT ( 1 / x ) dx ln v - ln ( 2 + v ) = - 4 ln x + 2 ln c.....Note the constant.... ln [ v / ( 1 + v ) ] = ln [ c^2 / x^2 ] v / ( 2 + v ) = c^2 / x^4 (y/x) / ( 2 + y/x ) = C / x^4 y / (2 x + y ) = C / x^4 x^4. y = C (2x + y ). ................General Solution.

OpenStudy (anonymous):

again I dont want the step by step I want to understand, no need to copy paste .